(20%) Data 1) For all calculations use following specifications A) Show the raw formula solving for the unknown variable B) Substitute into the raw formula with units C) Show final answer with units 2) For all calculations abide by the rules of significant fioures 1) (0%) Complete Table 1 below Microsoft's Equation Editor to show all work for each calculation with the Average Subject of the Tral Initial Final Change in Displacement Displacement (inches) Change in displacement (inches) displacemen (inches) t Test (inches) 22 0 2 0 24 3 0 19 4 0 26 5 0 22 Table 1
2) (10%) Show all work for specified calculations required to complete Table 1 A) (10%) Use the formula df= di+ Ad to show the calculation to compute the vertical displacement of the student for the first trial (first row of data) from Table 1 of Part 1 EquationObject Note: Edit the equation object to show the 3 lines, as outlined in the directions directly below the Data section. Note: Rearrange the raw formula to solve for Ad before substituting experimental values Show the calculation to compute the average vertical displacement of the student EquationObject Note: Edit the equation object to show the 3 lines directly below the Data section. B) (0% as outlined in the directions %) Conclusion: 1) For all calculations use Microsoft's Equation Editor to show all work for each calculation with the following specifications A) Show the raw formula solving for the unknown variable. B) Substitute into the raw formula with units. C) Show final answer with units 2) For all calculations abide by the rules of significant figures. 1) (10% Convert 6.0 feet to centimeters via the factor-label method using only the appropriate ratios 100 centimeter s 12 inches 5,280 feet 2.54 centimeter S 1 foot 1 mile linch 1meter from the following: 60 seconds 60 minute and 1 minute 1hour Show all work! Note: Though mathematical constants (such as 2 and t in the formula 27r) and many conversion factors (such as 60minutes/1hour or 60seconds/1minute or 12inches/1foot or 100cm/1m) precise. For example, 60minutes/1hour could be written as 60.000,000.../1.000,000..hour infinitely precise, measurements are only finitelv are Note: The conversion factor 2.54cm/in and the derived conversion factor 5,280ft/mile are accurate to the significant figures shown, these conversion factors are not infinitely precise. Hint: The answer is between 150cm and 200cm, inclusive Note: Only 1 line is required for a unit conversion EquationObject
Answer #1

1) Change in displacement = final displacement - initial displacement

     Average change in displacement = sum of all " change in displacement" / no of trials

Initial displacement final displacement change in displacement
0 22 22-0 = 22
0 24 24-0=24
0 19 19-0=19
0 26 26-0=26
0 22 22-0=22
Average change in displacement

22 241926+22 22.6inches 5

The LEAST number of significant figures in any number of the problem determines the number of significant figures in the answer.

By this rule rule answer should have 1 significant figure (because of '5' !).

Round off 22.6 to 1 significant figure

Answer = 20 inches